I have not had this in 30 years helping my 7th grade son in pre - algebra
can you solve and show steps please?
1 problem only,
5 3/7 B = - 7 13/28

Found 2 solutions by Alan3354, scott8148:
Answer by Alan3354(69443)

5 3/7 B = - 7 13/28
Change to fractions
38b/7 = -209/28
152b/28 = -209/28
152b = -209
b = -209/152
b = -11/8

Answer by scott8148(6628)

changing mixed numbers to fractions ___ 38/7 B = -209/28

multiplying by inverse to clear fraction ___ (7/38)38/7 B = -209/28 (7/38)

B = -(209 * 7) / (28 * 38)

reducing ___ B = -(11 * 1) / (4 * 2) = -11/8 ___ or ___ -1 3/8


we parents need to stick together